OpenStudy (anonymous):

Use the substitution method to evaluate the integral integrate dy/2(y^1/2)(1+y^1/2 )^2) dx, x=1..4

OpenStudy (wasiqss):

its easy jus write the equation using equations, im having trouble reading it

OpenStudy (wasiqss):

ohh i got it

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

962/15

OpenStudy (anonymous):

What's that Aron? I need steps if you can perform the intermediate steps

OpenStudy (anonymous):

Possible intermediate steps: integral 2 (1+sqrt(y))^2 sqrt(y) dy Factor out constants: = 2 integral (sqrt(y)+1)^2 sqrt(y) dy For the integrand (sqrt(y)+1)^2 sqrt(y), substitute u = sqrt(y) and du = 1/(2 sqrt(y)) dy: = 4 integral u^2 (u+1)^2 du Expanding the integrand u^2 (u+1)^2 gives u^4+2 u^3+u^2: = 4 integral (u^4+2 u^3+u^2) du Integrate the sum term by term and factor out constants: = 4 integral u^4 du+8 integral u^3 du+4 integral u^2 du The integral of u^2 is u^3/3: = 4 integral u^4 du+(4 u^3)/3+8 integral u^3 du The integral of u^3 is u^4/4: = 2 u^4+4 integral u^4 du+(4 u^3)/3 The integral of u^4 is u^5/5: = (4 u^5)/5+2 u^4+(4 u^3)/3+constant Substitute back for u = sqrt(y): = (4 y^(5/2))/5+(4 y^(3/2))/3+2 y^2+constant Now just substitute the required values!